### v1

# Hardware-Aware Quantum Kernel Design Based on Graph Neural Networks

Abstract
Quantum kernel methods have emerged as a promising direction in quantum machine learning (QML), offering a principled way to map classical data into high-dimensional quantum Hilbert spaces. While conceptually powerful, designing effective quantum kernels that adapt to both the target task and the constraints of near-term quantum hardware remains a nontrivial challenge. In this work, we propose HaQGNN, a hardware-aware quantum kernel design framework that integrates quantum device topology, noise characteristics, and graph neural networks (GNNs) to evaluate and select task-relevant quantum circuits. By predicting surrogate metrics related to fidelity and kernel performance, HaQGNN enables efficient circuit screening at scale. Feature selection is further incorporated to improve compatibility with limited-qubit systems and mitigate kernel degradation. Extensive experiments on three benchmark datasets, Credit Card, MNIST-5, and FMNIST-4, demonstrate that HaQGNN outperforms existing quantum kernel baselines in terms of classification accuracy. Our work highlights the potential of learning-based and hardware-aware strategies for advancing practical quantum kernel design on NISQ devices.
preprint: APS/123-QED
I Introduction
Quantum computing presents a compelling prospect for revolutionizing machine learning tasks by exploiting the fundamental properties of quantum mechanics, such as superposition, entanglement, and high-dimensional Hilbert spaces [1, 2]. Quantum machine learning (QML), as a important class of quantum algorithms tailored for the Noisy Intermediate-Scale Quantum (NISQ) era [3], has demonstrated practical applicability in diverse real-world tasks, including classification [4, 5, 6, 7, 8], generative modeling [9, 10, 11, 12], and learning physical systems [13, 14]. Among various QML approaches, quantum kernel methods have attracted increasing attention due to their solid theoretical foundation and interpretability [15, 16, 17]. By embedding classical data into quantum states through parameterized quantum circuits, quantum kernels define implicit feature maps capable of capturing complex nonlinear relationships beyond classical means.
Despite recent demonstrations of quantum kernel learning on real quantum devices, realizing practical quantum advantage remains challenging, especially under the constraints of Noisy Intermediate-Scale Quantum (NISQ) devices [18, 19, 20]. In particular, unsuitable quantum circuit designs may lead to vanishing kernel similarities [21], reduced expressivity, and degraded generalization [22, 23], severely limiting the performance of quantum kernels. These drawbacks are exacerbated when the data dimensionality is high or when the circuit depth and qubit count exceed the noise tolerance of near-term devices. To mitigate these limitations, several works have investigated circuit-aware and data-aware quantum kernel optimization. Some methods focus on tuning variational parameters in fixed circuit layouts [24, 25, 26], while others explore structural search via evolutionary or Bayesian algorithms [27, 28, 29]. However, most existing approaches either ignore hardware constraints such as qubit connectivity and gate noise, or suffer from excessive computational costs in searching large quantum circuit spaces.
Figure 1: An overview of handware-aware quantum kernel design.
To address these challenges, we propose a hardware-aware quantum kernel design framework based on Graph Neural Networks (HaQGNN), which efficiently integrates quantum hardware topology and device noise characteristics into the design of quantum feature maps, as illustrated in Fig. 1. The use of GNNs builds upon our prior work [30], where graph-based models were employed to predict specific properties of quantum circuits with high accuracy. Rather than formulating quantum kernel design as a joint optimization over circuit structure and parameters, our approach decouples the process by first generating hardware-aware candidate circuits and then employing graph neural networks to evaluate their potential performance. Leveraging the graph structure of quantum circuits, we employ GNNs to predict two surrogate metrics: the ”Probability of Successful Trials (PST)” [31], which reflects circuit fidelity under hardware noise, and kernel-target alignment (KTA), which correlates with downstream classification accuracy.
The HaQGNN framework consists of four steps: (1) candidate circuit generation based on hardware topology and native gate sets; (2) PST prediction via GNNs-1 to filter out low-fidelity circuits; (3) KTA prediction via GNNs-2 to select high-performing kernels; and (4) parameter optimization of the top-$k$ circuits for final task-specific performance. Furthermore, we incorporate feature selection to reduce input dimensionality, mitigating vanishing similarity issues and enhancing compatibility with NISQ devices. To the best of our knowledge, this is the first work to propose a GNN-based, hardware-aware framework that evaluates and selects quantum kernel circuits by jointly considering hardware topology, noise characteristics, and circuit structure, followed by task-specific parameter optimization. Extensive simulations on IBM quantum devices demonstrate the superiority of HaQGNN over several state-of-the-art quantum kernel baselines across multiple classification tasks and device configurations.
Our proposed HaQGNN framework offers several distinctive advantages that collectively enable efficient quantum kernel design for NISQ devices:
  • $\bullet$
    Hardware-awareness: HaQGNN explicitly incorporates real-device topology and noise information during both circuit generation and performance prediction, enhancing compatibility and circuit fidelity across different quantum hardware backends.
  • $\bullet$
    Two-stage GNN-based performance estimation: By employing two lightweight graph neural networks (GNNs-1 and GNNs-2), the framework accurately predicts fidelity (PST) and classification potential (KTA) of circuits with minimal simulation cost, achieving speedups of over 380× and 58000× compared to traditional metrics.
  • $\bullet$
    Noise-robust candidate selection: By filtering out low-fidelity circuits early in the pipeline, HaQGNN effectively mitigates the impact of gate errors and decoherence, leading to more reliable kernel performance on noisy devices.
  • $\bullet$
    Efficient support for high-dimensional input: The integration of feature selection techniques (e.g., mRMR) reduces data dimensionality while preserving discriminative information, helping avoid vanishing kernel similarity and enabling practical deployment on NISQ devices.
  • $\bullet$
    Superior classification accuracy: Extensive experiments on multiple benchmark tasks (e.g., Credit Card, MNIST-5, FMNIST-4) and across various IBM quantum devices (e.g., Perth, Lagos, Torino) demonstrate that HaQGNN consistently outperforms several state-of-the-art quantum kernel baselines.
The remainder of this paper is organized as follows. Section II introduces background knowledge on quantum computing, kernel methods, and graph neural networks. Section III describes the proposed HaQGNN framework in detail. Section IV presents numerical results and comparative evaluations. Finally, Section V concludes the paper and discusses future research directions.
II Background
II.1 Quantum computation fundamental
A quantum bit (qubit) can exist in a linear superposition of the two computational basis states, $\left|0\right\rangle$ and $\left|1\right\rangle$, in contrast to a classical bit that can only be in a definite state of either 0 or 1. The state of a single qubit is generally represented as $\left|\psi\right\rangle=\alpha\left|0\right\rangle+\beta\left|1\right\rangle$, where $\alpha$ and $\beta$ are complex amplitudes satisfying the normalization condition $\left|\alpha\right|^{2}+\left|\beta\right|^{2}=1$. For a system of multiple qubits, the overall quantum state is represented by the tensor product of individual qubit states, enabling the description of entangled states that cannot be factorized into separate 1-qubit components. Quantum gate, as a unitary matrix, transforms the quantum state to another and preserves the norm of the quantum state. Any unitary can be decomposed into a set of basis gates. The basis gate sets explored in this study are $S=\left\{R_{x},R_{z},CZ,I\right\}$ and $S^{\ast}=\left\{R_{z},X,CNOT,I\right\}$, which primarily consist of the 1-qubit gates $I$, $X$, $R_{x}$ and $R_{z}$, whose matrix representations are given as follows:
$I=\begin{pmatrix}1&&0\\ 0&&1\end{pmatrix}$ $X=\begin{pmatrix}0&&1\\ 1&&0\end{pmatrix}$ (1)
$R_{x}\left(\theta\right)=\begin{pmatrix}\cos\frac{\theta}{2}&-i\sin\frac{\theta}{2}\\ -i\sin\frac{\theta}{2}&\cos\frac{\theta}{2}\end{pmatrix}$ $R_{z}\left(\theta\right)=\begin{pmatrix}e^{-i\frac{\theta}{2}}&0\\ 0&e^{i\frac{\theta}{2}}\end{pmatrix}$ (2)
In addition, the 2-qubit gates $CNOT$ and $CZ$ are also included, with their corresponding matrix representations defined as:
$CNOT=\begin{pmatrix}1&&0&&0&&0\\ 0&&1&&0&&0\\ 0&&0&&0&&1\\ 0&&0&&1&&0\end{pmatrix}$ $CZ=\begin{pmatrix}1&&0&&0&0\\ 0&&1&&0&0\\ 0&&0&&1&0\\ 0&&0&&0&-1\end{pmatrix}$ (3)
II.2 Mechanism of quantum kernel
Kernel methods provide a powerful theoretical framework for performing nonlinear and nonparametric learning tasks, owing to their generality and strong interpretability [23]. Assume that both the training and test samples are drawn from the same input-output space $\mathcal{X}\times\mathcal{Y}$. The training dataset is denoted as $\mathcal{D}=\left\{\bm{x}_{i},y_{i}\right\}_{i=1}^{n}\subset\mathcal{X}\times\mathcal{Y}$, where $\bm{x}_{i}$ represents the input features of the $i$-th sample and $y_{i}$ denotes the corresponding label. The fundamental idea of traditional kernel methods is to map input data $\bm{x}_{i}$ from $\mathbb{R}^{d}$ to a higher-dimensional feature space $\mathbb{R}^{q}$, i.e., $\phi\left(\cdot\right):\mathbb{R}^{d}\to\mathbb{R}^{q}$ with $q\gg d$, such that data that are not linearly separable in the original space may become linearly separable in the transformed space. However, when $q$ is large, explicitly computing the mapped feature vectors $\phi\left(\bm{x}_{i}\right)$ becomes computationally expensive. To address this issue, kernel methods leverage the kernel trick, which enables implicit computation by constructing a kernel matrix $K\in\mathbb{R}^{n\times n}$, where each element is defined as the inner product of feature vectors in the high-dimensional space:
$\left[K\right]_{i,j}=\left[K\right]_{j,i}=\left\langle\phi\left(\bm{x}_{i}\right),\phi\left(\bm{x}_{j}\right)\right\rangle,\forall i,j\in\left[n\right]$ (4)
This approach enables efficient similarity computation with a time complexity of $\mathcal{O}\left(d\right)$, without explicitly performing the high-dimensional mapping. The aim of kernel learning is to infer a hypothesis function $h\left(\bm{x}_{i}\right)=\left\langle\bm{w}^{\ast},\phi\left(\bm{x}_{i}\right)\right\rangle$ with high test accuracy by utilizing the constructed kernel matrix $K$. The optimal parameter $\bm{w}^{\ast}$ is obtained by minimizing the following loss function:
$\mathcal{L}=\lambda\left\langle\bm{w},\bm{w}\right\rangle+\sum_{i=1}^{n}\left(\left\langle\bm{w},\phi\left(\bm{x}_{i}\right)\right\rangle\right)^{2}$ (5)
The performance of kernel methods heavily depends on the choice of the feature mapping function $\phi\left(\cdot\right)$, or equivalently, the kernel function $\kappa\left(\cdot,\cdot\right)=\left\langle\phi\left(\cdot\right),\phi\left(\cdot\right)\right\rangle$. Various kernel functions have been proposed to accommodate different task scenarios, such as the radial basis function (RBF) kernel, Gaussian kernel, circular kernel, and quantum kernel.
Compared to classical kernel methods, the primary distinction of quantum kernel approaches lies in the implementation of the feature mapping function $\phi\left(\cdot\right)$. For an $N$-qubit quantum kernel model, the feature mapping of a sample $\bm{x}_{i}$ is realized through a parameterized quantum circuit $U_{c}\left(\bm{x}_{i},\bm{\theta}\right)$ applied to the initial quantum state $\left|0\right\rangle^{\otimes N}$, resulting in the quantum state $\left|\varphi\left(\bm{x}_{i},\bm{\theta}\right)\right\rangle=U_{c}\left(\bm{x}_{i},\bm{\theta}\right)\left|0\right\rangle^{\otimes N}$. Here, $\bm{\theta}$ denotes trainable parameters and the corresponding density matrix is denoted as $\rho\left(\bm{x}_{i},\bm{\theta}\right)=\left|\varphi\left(\bm{x}_{i},\bm{\theta}\right)\right\rangle\left\langle\varphi\left(\bm{x}_{i},\bm{\theta}\right)\right|$. The objective of quantum kernel learning is to utilize the quantum kernel matrix $W\left(\bm{x},\bm{\theta}\right)$, defined as
$\left[W\left(\bm{x},\bm{\theta}\right)\right]_{i,j}=Tr\left(\rho\left(\bm{x}_{i},\bm{\theta}\right)\rho\left(\bm{x}_{j},\bm{\theta}\right)\right),\forall i,j\in\left[n\right]$ (6)
to infer a hypothesis as formulated in Eq. 5, with the goal of achieving higher test accuracy compared to classical kernel methods. Therefore, the performance of a quantum kernel largely depends on the design of the parameterized circuit $U_{c}\left(\cdot,\bm{\theta}\right)$. It is worth noting that the implementation of $U_{c}\left(\cdot,\bm{\theta}\right)$ is flexible and can take various forms, such as circuits constructed with predefined architectures, hardware efficient circuits, or circuits randomly generated from a specified quantum gate set.
II.3 Graph neural networks
Graph Neural Networks (GNNs) have become a powerful framework for learning over graph-structured data, where relationships among entities play a critical role in the underlying data distribution [32]. By iteratively aggregating and transforming information from neighboring nodes through message-passing mechanisms, GNNs are capable of capturing both local topology and global structural dependencies [33]. This capacity to model complex relational patterns has led to their widespread adoption across a variety of domains, including social network analysis [34], molecular property prediction [35, 36, 37], and recommender systems [38, 39]. Notably, variants such as Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), and GraphSAGE have significantly advanced the state-of-the-art in graph representation learning tasks [40, 41, 42, 43, 44]. In the context of quantum computing, quantum circuits can be naturally represented as graphs, where qubits correspond to nodes and quantum gates or couplings define the edges. This structural analogy provides a compelling motivation for applying GNNs to quantum systems, enabling efficient learning and prediction tasks grounded in the intrinsic connectivity of quantum architectures.
III Method
In this section, we present the technical details of our proposed hardware-aware quantum kernel design framework. An overview of the framework is illustrated in Fig. 1, which consists of four main steps:
Step 1: Based on the hardware topology and noise characteristics of the quantum device, a set of qubits is selected such that they exhibit favorable connectivity and minimal susceptibility to noise. A large number of candidate quantum circuits are then generated on the selected qubits, following the supported native gate set and qubit connectivity of the target device.
Step 2: A graph neural network (GNNs-1) is employed to predict the ”Probability of Successful Trials (PST)”, a metric correlated with circuit fidelity. This step enables the early rejection of low-fidelity candidate circuits, thereby filtering out circuits with potentially higher fidelity.
Step 3: A second graph neural network (GNNs-2) is designed to estimate the kernel-target alignment (KTA), a surrogate metric for classification accuracy, over the high-fidelity circuits identified in Step 2. The top-$k$ circuits with the highest predicted KTA values are selected for further refinement.
Step 4: A classical optimizer is employed to train the parameters of the top-$k$ high-KTA circuits with the high PST values from scratch, and the circuit with the best task-specific classification accuracy is selected as the final quantum kernel circuit.
We divide the hardware-aware quantum kernel design into three parts: candidate circuit generation (Step 1, Section III.1), metric prediction based on graph neural networks (Step 2-3, Section III.2), circuit training and selection (Step 4, Section III.3).
III.1 Candidate circuit generation
The process of generating candidate quantum circuits is primarily divided into three stages: 1) A specific number of qubits with preserved topological connectivity are selected from the quantum device. This selection process takes into account the noise characteristics of individual qubits and their couplings, aiming to identify a subgraph of the device topology with the lowest average noise. Such the selection strategy improves hardware efficiency and enhances the fidelity of the generated circuits. 2) Candidate circuits are generated based on the selected qubit and the native gate set supported by the quantum device. During the circuit generation process, each inserted quantum gate must act on qubits that are physically connected on the hardware. This ensures that qubit mapping is implicitly achieved during circuit construction, thereby avoiding the insertion of costly SWAP gates during the compilation phase. As a result, the overall circuit depth is reduced, and noise accumulation is mitigated. 3) In the candidate quantum circuits, parameterized quantum gates are selected at different positions to enable flexible data embedding, which helps improve the performance of the quantum kernel.
III.1.1 Subgraph selection in device topology
Due to the non-negligible differences in noise levels across individual qubits and qubit couplings within quantum devices, we perform a selection process to identify a set of $N$ qubits that are less affected by noise. This selection aims to improve hardware efficiency and maintain higher circuit fidelity. Fig. 2 illustrates examples of selected subgraphs with $N=4,5,7$ and $8$ extracted from the 133-qubit IBM Torino topology.
Figure 2: Subgraphs with $N=4,5,7$ and $8$ selected from the 133-qubit IBM Torino topology.
Selecting a subgraph with the desired $N$ qubits requires a comprehensive consideration of several hardware-specific factors in the quantum device topology, including 1-qubit gate errors, readout errors, 2-qubit gate errors between coupled qubits, and the degree of connectivity among qubits. The target is to identify a subgraph with high gate fidelity and high qubit connectivity, while maintaining low readout errors. By incorporating device-specific noise information, the subgraph selection strategy ensures that the subsequently generated quantum circuits are noise-aware, with detailed procedures provided in Appendix A.
III.1.2 Generating candidate circuits based on the subgraph
After selecting a high-fidelity subgraph, denoted as $G$, we generate $M$ candidate quantum circuits based on the structure of $G$ and the set of quantum gates supported by the device, denoted as $S$. This process ensures that the generated circuits are hardware-aware. For example, as shown in Step 1 of Fig. 1, candidate circuits are generated from the subgraph of the IBM Torino using the gate set $S=\left\{R_{x},R_{z},CZ,I\right\}$. The detailed procedure for candidate circuit generation is provided in Algorithm III.1.2. During the candidate circuit generation process, we guide the selection of 2-qubit gate between different coupled qubits based on their gate errors. Specifically, the probability of inserting a 2-qubit gate between qubits $q_{i}$ and $q_{j}$, denoted as $P\left(q_{i},q_{j}\mid e_{j}\right)$, is defined as follows:
$\begin{array}[]{c}P\left(q_{i},q_{j}\mid e_{j}\right)=\frac{1/e_{j}}{{\textstyle\sum_{j}1/e_{j}}}\end{array}$ (7)
where the 2-qubit gate error $e_{j}$ corresponds to the edge $q_{i}q_{j}$ in subgraph $G$. 1-qubit gate errors are not considered because 2-qubit gate errors are typically an order of magnitude higher and have a more significant impact on circuit performance. Moreover, since the construction of candidate circuits is guided by noise information, the resulting circuits are inherently noise-aware. By repeatedly executing the procedure in Algorithm III.1.2, a large number of hardware- and noise-aware candidate quantum circuits can be generated, laying the foundation for subsequent GNN-based training and the identification of optimal quantum circuits.
Algorithm 1: Generate candidate circuit
1:Device subgraph $G\left(\left\{V_{i}\right\},\left\{E_{j}\right\}\right)$; 2-qubit gate error $\left\{e_{j}\right\}$; The gate set $S$; The number of gates $P$.
2:Candidate circuit $C$.
3:for $k$ in range($P$do
4:     Sample the gate $g_{k}$ from $S$ on average $\triangleright$ Sample a gate to be added, e.g., $R_{x}$, $R_{z}$, or $CZ$.
5:     if $g_{k}$ is a 1-qubit gate then
6:         Select a qubit $q_{i}\in\left\{V_{i}\right\}$ with average probability from subgraph $G$triangleright Select a qubit from the subgraph as the target qubit of the quantum gate.